OpenStudy (anonymous):

X^2-9/x+2 divided by x-3/x-2 ?

OpenStudy (anonymous):

i would suggest factoring the numertaor of the first fraction

OpenStudy (anonymous):

your eq is (x-3)(x+3)/(x+2) / (x-3)/(x-2) ---> (x+3)(x-2)/(x+2)

OpenStudy (anonymous):

if you've got (a/b) / (c/d) you can change it into (a/b)*(d/c) i think this will make it easier

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

first invert. but don't multiply. factor and cancel

OpenStudy (anonymous):

\[\frac{x^2-9}{x+2}\times \frac{x-2}{x+3}\] factor \[x^2-9=(x+3)(x-3)\] and cancel

OpenStudy (anonymous):

I have the correct answer as \[(x-2)(x+3)\div x+2\] but i'm trying to figure out the reasoning behind it

OpenStudy (anonymous):

thats what i told you if you've got (a/b)/(c/d) its the same as (a*c)/(b*d) and from there u just look at satellites answer